OpenStudy (anonymous):

given the points A(9, 11) and B(-3, -5). Find AB and M, The Midpoint Of AB . please help

OpenStudy (accessdenied):

\[ D = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2} \\ \\ M = \left(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\right) \\~\\ \quad \large{\text{Given } (x_1,y_1) \text{ and } (x_2,y_2).} \]

OpenStudy (anonymous):

sorry but i dont get it :(

OpenStudy (accessdenied):

Basically you just plug in the values into the formula and evaluate. Or did you not learn these formulas? There are other ways to find it using Pythagorean Theorem, if you're familiar with that...